OpenStudy (anonymous):

Write the explicit formula for the geometric sequence. a1 = -5 a2 = 20 a3 = -80 A. an = -5 • (-4)n B. an = -5(-4)n-1 C. an = -4(-5)n-1 D. an = -5 • (4)n

OpenStudy (jmark):

In geometric sequence, the formula to find the nth term is given below a_n = ar^(n-1), where a = first term, r = common difference, n = total number of term. Note: common difference can be found by dividing second term by the first term. Since a_1 = -5, a_2 = 20 and a_3 = -80 We are given with three terms so n = 3 r = a_2/a_1 = 20/-5 = -4 a_n = -5(-4)^2 CBSE Class 6 Math Question Paper - http://education-help-online.blogspot.in/2014/07/cbse-class-6-math-question-paper.html